write an exponential function of the form y=ab^x that includes the given points..

(-1,3/8),(1,3/2)

thanks for the help

Answer by stanbon(75887)

write an exponential function of the form y=ab^x that includes the given points..
(-1,3/8),(1,3/2)
------------------------------
Substitute for x and y to solve for a and b:
(3/8) = ab^(-1)
(3/2) = ab^(1)
---------------------
Divide the 2nd equation by the 1st to solve for "b":
4 = b^2
b = 2
----
Now solve for "a":
(3/2) = a(2)^1
a = 3/4
--------------------
Equation:
y = (3/4)(2)^x
-----
Cheers,
Stan H.